OpenStudy (anonymous):

Using point-slope form, write the equation of the line that passes through (2, –1) and has a slope of –5. A. (y – 1) = –5(x + 2) B. (y – 2) = –5(x – 1) C. (y + 1) = –5(x – 2) D. (y + 2) = –5(x – 1)

OpenStudy (harsimran_hs4):

given a point (a, b) and slope m eq of line is y-b = m (x-a) can you proceed now ?

OpenStudy (anonymous):

ill try thanks

OpenStudy (anonymous):

do you know how to plug in the points and the slope?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

well you can see that there is y1 and x1 and m right?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

it is similar to the x and the y and m is the slope

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

y1 = y and x1 = x m = slope

OpenStudy (anonymous):

are you still confused?

OpenStudy (anonymous):

I will help you

OpenStudy (anonymous):

yeah :/

OpenStudy (anonymous):

well you know that you have the coordinate (x,y) right?

OpenStudy (anonymous):

yeah ik that

OpenStudy (anonymous):

(2, –1) well that means that the x and the y is actually the x1 and the y1 (x1,y1) so when you plug in the points you can get the answer easily

OpenStudy (anonymous):

alright ill try it thanks

OpenStudy (anonymous):

and btw the equation is y - y1 = m(x- x1)

OpenStudy (anonymous):

i think its A the answer

OpenStudy (anonymous):

no quite it is the last one

OpenStudy (anonymous):

D ?

OpenStudy (anonymous):

I mean second

OpenStudy (anonymous):

to the last one

OpenStudy (anonymous):

C

OpenStudy (anonymous):

thank you for helping :)

OpenStudy (anonymous):

yw